Please show ALL steps for each partSuppose that, in a certain population, heights of males are normally distributed with mean M1 = 178cm and standard deviation
Answer #1

x= height of male Y= height of Female Given that X n Normal (198, 4²) Yn Normal (170, 3²) We know that - X., X2, Yz oo. No nNTherefore, je in Normal (178, 4 ) û = 1782 M = 178 San And also a random sample of as females ase Selected ther their Average(a) . we have tn Normal (178, 0.8²) ī n Normal (170, 0.6) X = Avg. height of mate : Ta Aug, height of female The Pobability t(6) The Probability that the Average height of the females is between 169 cm and 171 con is P(169 <ř <17) = P(169-4 < z < 171c) The Probability that the average height of the males is more than 10cm greater than the Average height of the females is LHere X, X, Xã - - - x = 4 ( 179 ) - Y,,Year. Yes NN (170,32) Lek K x, + X₂ 4 - .Xast Y, A Yet - Y25 50 F(x) = 25 (118) +25 (1P(1726 x< 175) = P(-4ę732) ap(222) - P(z sah) -0.97725 - 0.00003 P (172<x< 195) = 0.97722! CS Scanned with CamScanner